# Hermite trace polynomials and chaos decompositions for the Hermitian Brownian motion

### v2

@@ 5.1.3. Polynomial subalgebra
Denote $X(h)=\mathnormal{T}\left({(01)}\otimes h\right)$. Then
$X(h_{1})\ldots X(h_{n})=\mathnormal{T}\left({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right).$
Since any two long cycles are conjugate,
and is a unital star-subalgebra of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$. It is not closed under contractions or conditional expectations. In particular, the corresponding $\mathnormal{W}\left({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)$ operators do not in general belong to this subalgebra.
@@
We will use induction on $n$. For $n=0$, $\alpha=(0)$. Suppose that for each $\beta\in S_{0}(k)$, $k<n$, $\mathnormal{T}\left(\beta\otimes_{s}F\right)$ is in the algebra generated by the conditional expectations of elements of $\mathcal{P}(\mathcal{H}_{\mathbb{R}})$. It suffices to show that $\mathnormal{T}\left({(0)(1\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)$ is in it. Let $h\perp\mathcal{H}^{\prime}_{\mathbb{R}}=\mathrm{Span}\left(h_{1},\ldots,h_{n}\right)$ be a non-zero vector. Then
@@
$\begin{split}&\varphi\left[X(h)\mathnormal{T}\left({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)X(h)\ |\ \mathcal{H}^{\prime}\right]\\ &\quad\varphi\left[X(h)(\mathnormal{W}\left({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)+\text{ lower order terms})X(h)\ |\ \mathcal{H}^{\prime}\right]\\ &\quad=\left\|h\right\|^{2}\mathnormal{W}\left({(0)(1\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)+\text{ lower order terms}.\end{split}$
The result follows. ∎
@@
On $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ it is implemented by
@@
In its representation on $\mathcal{F}_{q}(\mathcal{H})$, for $h\in\mathcal{H}_{\mathbb{R}}$, $X(h)$ is essentially self-adjoint.
@@
Clearly $X({h})$ is symmetric. So it suffices to show that for each $\alpha\in S_{0}(k)$, $\mathnormal{T}\left(\alpha\otimes F\right){(0)}$ is an analytic vector for it. Indeed,
$\begin{split}&\frac{1}{n}\left\|X(h)^{n}\mathnormal{T}\left(\alpha\otimes F\right){(0)}\right\|^{1/n}\\ &\quad=\frac{1}{n}\varphi\left[\mathnormal{T}\left({\alpha^{-1}\cup(01\ldots n)\cup\alpha}\otimes(\bar{F}\otimes h^{\otimes 2n}\otimes F)\right)\right]^{1/2n}\\ &\quad=\frac{1}{n}\left(\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n+2k)}q^{(n+k)-\operatorname{\mathrm{cyc}}_{0}(\pi(\alpha^{-1}\cup(01\ldots n)\cup\alpha))}C_{\pi}(\bar{F}\otimes h^{\otimes 2n}\otimes F)\right)^{1/2n}\\ &\quad\leq\frac{1}{n}\left(\frac{(2n+2k)!}{2^{n+k}(n+k)!}\left\|F\right\|^{2}\left\|h\right\|^{2n}\right)^{1/2n}\\ &\quad\sim\frac{1}{n}2^{1/2}(n+k)^{1/2}e^{-1/2}\left\|h\right\|\rightarrow 0.\qed\end{split}$
@@
$\begin{split}a^{-}_{(01)}(h)(\alpha\otimes F)&=q\sum_{k\neq\alpha^{-1}(0)}P^{[0,n]\setminus\left\{k\right\}}_{[0,n-1]}((0k)\alpha)|_{\left\{k\right\}^{c}}\otimes r^{-}_{k}(h)F\\ &\qquad+P^{[0,n]\setminus\left\{\alpha^{-1}(0)\right\}}_{[0,n-1]}\alpha|_{\left\{\alpha^{-1}(0)\right\}^{c}}\otimes r^{-}_{\alpha^{-1}(0)}(h)F,\end{split}$
where if $\alpha(0)=0$, the final term is absent.
The distribution of $\mathnormal{W}\left({(01)}\otimes h\right)$ is the unnormalized average empirical distribution of a GUE matrix with mean 00 and variance $\left\|h\right\|$.
@@
$\begin{split}&\mathnormal{W}\left(\alpha\otimes F\right)\ \mathnormal{W}\left((01)\otimes h\right)\\ &\quad=\mathnormal{W}\left((0\ n+1)\alpha\otimes r^{+}(h)F\right)\\ &\quad\quad+\sum_{k=1}^{n}q^{\operatorname{\mathrm{cyc}}_{0}(((0k)\alpha)|_{\left\{k\right\}^{c}})-\operatorname{\mathrm{cyc}}_{0}((0k)\alpha)+1}\mathnormal{W}\left(P^{[0,n]\setminus\left\{k\right\}}_{[0,n-1]}((0k)\alpha)|_{\left\{k\right\}^{c}}\otimes r^{-}_{k}(h)F\right)\\ &\quad=\mathnormal{W}\left((0\ n+1)\alpha\otimes r^{+}(h)F\right)\\ &\quad\quad+q\sum_{k\neq\alpha^{-1}(0)}\mathnormal{W}\left(P^{[0,n]\setminus\left\{k\right\}}_{[0,n-1]}((0k)\alpha)|_{\left\{k\right\}^{c}}\otimes r^{-}_{k}(h)F\right)\\ &\quad\quad+\mathnormal{W}\left(P^{[0,n]\setminus\left\{\alpha^{-1}(0)\right\}}_{[0,n-1]}\alpha|_{\left\{\alpha^{-1}(0)\right\}^{c}}\otimes r^{-}_{\alpha^{-1}(0)}(h)F\right).\end{split}$
@@
which should be compared with Theorem 22.12 in [NS06]. ∎
@@ 5.2. The relation to a construction by Bożejko and Guţă
We contrast the algebra $\mathcal{P}(\mathcal{H}_{\mathbb{R}})$ with a construction from [BG02]. In section 5 of that paper, Bożejko and Guţă considered the Fock space with the inner product
@@
and the corresponding expression is zero for an odd number of factors (compare with equation (10)). Here $c(\pi)$ is again the number of cycles of a permutation corresponding to a partition $\pi$, but this correspondence is more subtle. For a pair partition $\pi$, there is a unique non-crossing partition $\tilde{\pi}$ with the same openers and closers as $\pi$. If $i\stackrel{{\scriptstyle\pi}}{{\sim}}j$ and $i\stackrel{{\scriptstyle\tilde{\pi}}}{{\sim}}k$, then for the corresponding permutation $\sigma$, $\sigma(i)=j$ if $i<j$, and $\sigma(i)=k$ if $k<i$ (it is easy to check that this is an alternative). In other words, $\sigma$ is a permutation with an upper partition $\pi$ and the lower partition $\tilde{\pi}$ in the sense of Corteel [Cor07]. Then $c(\pi)$ is the number of cycles of $\sigma$.
@@
Combining our construction with [BG02], one could consider a Fock space with the inner product
@@
Here the coefficient of $\beta$ is the generating function of the number of primitive factorizations of $\beta$. See [MN10].
@@ 6.1. Background
Let $\left\{x_{i}:i\in S\right\}$ be a collection of non-commuting variables. Informally, a trace polynomial in these variables is a polynomial in these variables as well as in a tracial functional $\operatorname{\mathrm{tr}}$ applied to these variables. See [Céb13] for a formal definition using a universal property.
For the purposes of this paper, we will use a more constructive definition. Let $\alpha\in S_{0}(n)$. Denote
$\operatorname{\mathrm{tr}}_{\alpha}[x_{1},\ldots,x_{n}]=\prod_{i\text{ in the cycle starting with $0$}}x_{i}\prod_{\text{other cycles}}\operatorname{\mathrm{tr}}\left[\prod_{i\text{ in the cycle}}x_{i}\right].$
For example, for $\alpha=(024)(137)(56)$,
$\operatorname{\mathrm{tr}}_{\alpha}[x_{1},\ldots,x_{7}]=x_{2}x_{4}\operatorname{\mathrm{tr}}[x_{1}x_{3}x_{7}]\operatorname{\mathrm{tr}}[x_{5}x_{6}].$
Compare with Notation 22.29 in [NS06]. Then $\operatorname{\mathrm{tr}}_{\alpha}[x_{1},\ldots,x_{n}]$ is a trace monomial, and a trace polynomial is a linear combination of trace monomials.
Next, let $\mathcal{A}$ be a unital algebra, $\mathcal{C}$ its center, and $F:\mathcal{A}\rightarrow\mathcal{C}$ a unital $\mathcal{C}$-bimodule linear map. For any $\alpha\in S_{0}(n)$, we can similarly form $F_{\alpha}(a_{1},\ldots,a_{n})\in\mathcal{A}$, and consider it as the application of the trace monomial $\operatorname{\mathrm{tr}}_{\alpha}[x_{1},\ldots,x_{n}]$ to the elements $a_{1},\ldots,a_{n}$. See [Céb13].
We extend the notation to $F_{\eta}$ for $\eta\in\mathbb{C}[S_{0}(n)]$ by linearity.
@@ 6.1.1. Invariant theory of $N\times N$ matrices
where $\operatorname{\mathrm{Tr}}$ is the (un-normalized) trace on $M_{N}(\mathbb{C})$. Indeed, since for the purposes of this expansion, $x_{ij}^{(k)}$ and $(x_{ij}^{(k)})^{\ast}$ can be considered as independent variables, this follows directly from the first Procesi-Razmyslov theorem [Pro76]. The result is usually formulated using $GL(n)$-invariance. However, the argument ultimately reduces to Schur-Weyl duality, for which (in the case of inner product spaces) unitary invariance is sufficient.
@@ 6.1.2. Hermitian Brownian motion
Let $\left\{b_{ij}(h):h\in\mathcal{H}_{\mathbb{R}}\right\}$ be $N^{2}$ standard independent Gaussian processes indexed by the same real Hilbert space, represented on the same probability space. Define the $N\times N$ Hermitian Gaussian process $\left\{X(h):h\in\mathcal{H}_{\mathbb{R}}\right\}$ by
$X(h)_{ij}=\begin{cases}\frac{1}{\sqrt{2N}}(b_{ij}(h)+\sqrt{-1}b_{ji}(h)),&i<j,\\ \frac{1}{\sqrt{N}}b_{ij}(h),&i=j,\\ \frac{1}{\sqrt{2N}}(b_{ij}(h)-\sqrt{-1}b_{ji}(h)),&i>j.\\ \end{cases}$
@@
Denote
@@ 6.2. The isomorphisms
Numerous corollaries follow by combining Theorem 6.4 with results from earlier in the article. We only list a few of them explicitly. Others include Proposition 4.8 (chaos decomposition in the univariate case), 5.6 (expansion of the Hermite polynomial and stochastic integral), 5.10 (product formula for Hermite polynomials and stochastic integrals), 5.16 (linearization coefficients), and Corollary 5.11 (Factorization of the Hermite polynomial for orthogonal arguments).
@@
Let $\mathcal{H}=L^{2}(\mathbb{R}_{+},dx)$. Then for each $\alpha$, $H_{\alpha}(X(\mathbf{1}_{[0,t]}))$ is a martingale.
For Hermite polynomials of matrix argument, the martingale property was proved in [Law08] by generating function methods. See Section 6.2.1 for the connection.
@@
In the case $\mathcal{H}_{\mathbb{R}}=L^{2}(\mathbb{R}_{+},dx)$ and $F\in L^{2}(\mathbb{R}_{+}^{n},dx^{\otimes n})$, we may identify $\mathcal{E}[\mathnormal{W}\left(\eta\otimes_{s}F\right)]$ with a stochastic integral
@@
$\mathcal{E}[\mathnormal{W}\left(\eta\otimes(\mathbf{1}_{J_{1}}\otimes\ldots\otimes\mathbf{1}_{J_{n}})\right)]=\operatorname{\mathrm{Tr}}_{\eta}[X(\mathbf{1}_{J_{1}}),\ldots,X(\mathbf{1}_{J_{n}})]\$
@@
[…]
$(:\!\!Y\!\!:)_{ij}=:\!\!(Y_{ij})\!\!:$
is also equivariant.
@@
It suffices to prove the result when each $Y_{ij}$ is a homogeneous polynomial of degree $n$. Let $Y=P(X(h):h\in\mathcal{K})$ and $:\!\!Y\!\!:=Q(X(h):h\in\mathcal{K})$. Let $U\in U_{N}(\mathbb{C})$. Since each $:(Y_{ij}):$ is orthogonal to $\mathcal{P}_{n-1}(\mathcal{K})$, for any such polynomial $R$,
@@
The matrix with entries $:\!\!(\operatorname{\mathrm{Tr}}_{\alpha}[X(h_{1}),\ldots,X(h_{n})])_{\ell r}\!\!:$ has leading term $\operatorname{\mathrm{Tr}}_{\alpha}[X(h_{1}),\ldots,X(h_{n})]$, polynomial entries, and (by Lemma 6.10) is equivariant. Therefore it is of the form
@@
In [Bia97], Biane defined the matricial Segal-Bargmann transform. We outline the construction, and send the reader to the original article for details. The ordinary Segal-Bargmann transform is defined in Section 2.4. Apply that construction with $\mathcal{H}_{\mathbb{R}}=M_{N}^{sa}(\mathbb{C})$, the space of Hermitian matrices. Upgrade the Segal-Bargmann transform to its matricial version by applying it entry-by-entry. Then it follows from Theorem 6.11 that, as in Lemma 7 and Theorem 7 of [Bia97],
@@
$h_{\lambda}(x_{1},\ldots,x_{N})=\mathcal{E}[\mathnormal{W}\left(\lambda\otimes h^{\otimes n}\right)].$
@@
For $\lambda\in\operatorname{\mathrm{Par}}(n)$, the Hermite polynomial of matrix argument (for $\beta=2$) is the symmetric polynomial in $\left\{x_{1},\ldots,x_{N}\right\}$ with leading term $\frac{\left|\lambda\right|!}{c_{\lambda}}s_{\lambda}$ which is an eigenfunction of the operator $D^{\ast}-E^{\ast}$ with eigenvalue $-n$ (note a misprint in [DES07]). Here (see Corollary 7.1.7.4 in [Sta99])
@@
is the Schur polynomial and $c_{\lambda}=\prod(\lambda_{i}+\lambda_{j}^{\prime}-i-j+1)$ is the hook length. See [BF97, DES07, For10] for more details.

### v3

@@ 5.1.3. Polynomial subalgebra
Definition 5.30.
The polynomial subalgebra of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ is
@@
Denote $X(h)=\mathnormal{T}\left({(01)}\otimes h\right)$, so that $\mathcal{P}(\mathcal{H}_{\mathbb{R}})$ is generated by these operators explaining the name. Also,
$X(h_{1})\ldots X(h_{n})=\mathnormal{T}\left({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)$
and any other long cycle is conjugate to the standard one, so the two expressions in the preceding definition are equivalent. $\mathcal{P}(\mathcal{H}_{\mathbb{R}})$ is a unital star-subalgebra of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$. It is not closed under contractions or conditional expectations. In particular, the operators $\mathnormal{W}\left({(01\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)$ do not in general belong to this subalgebra.
@@
By the preceding remark, it suffices to show that the algebra generated by conditional expectations contains the center $\mathrm{Span}\left(\mathnormal{T}\left(\alpha\otimes_{s}F\right):\alpha(0)=0\right)$. By Definition 5.8 and linearity, the center is generated by $\mathnormal{T}\left(\beta\otimes_{s}F\right)$, where $\beta\in S(n)$ is a single cycle not containing 0, and $F$ is a simple tensor. We will use induction on $n$. Suppose that for each $\beta\in S(k)$, $k<n$, $\mathnormal{T}\left(\beta\otimes_{s}F\right)$ is in the algebra generated by the conditional expectations of elements of $\mathcal{P}(\mathcal{H}_{\mathbb{R}})$. It then suffices to show that $\mathnormal{T}\left({(0)(1\ldots n)}\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)$ is in it. Let $h\perp\mathcal{H}^{\prime}_{\mathbb{R}}=\mathrm{Span}\left(h_{1},\ldots,h_{n}\right)$ be a non-zero vector. Then
@@
[…]
The result follows by induction. ∎
@@
Here for $\alpha\in S_{0}(n)$, the permutation $\alpha|_{\left\{0\right\}^{c}}$ is obtained by removing 0 from its cycle in $\alpha$, similarly to Notation 3.6. On $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ this conditional expectation is implemented by
@@
For $q\in\mathcal{Z}$, in its representation on $\mathcal{F}_{q}(\mathcal{H})$, for $h\in\mathcal{H}_{\mathbb{R}}$, $X(h)$ is essentially self-adjoint.
@@
Clearly X(h)X({h}) is symmetric. So by Nelson’s analytic vector theorem, it suffices to show that its domain contains a subset of analytic vectors which is total (i.e. their span is dense). We verify that for each $\alpha\in S_{0}(k)$, $\mathnormal{T}\left(\alpha\otimes F\right){(0)}$ is an analytic vector for it. Indeed, using the relation in Theorem 5.15 between the inner product on $\mathcal{F}_{q}(\mathcal{H})$ and the state on $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$,
$\begin{split}&\frac{1}{n}\left\|X(h)^{n}\mathnormal{T}\left(\alpha\otimes F\right){(0)}\right\|_{q}^{1/n}\\ &\quad=\frac{1}{n}\varphi\left[\mathnormal{T}\left({\alpha^{-1}\cup(01\ldots n)\cup\alpha}\otimes(\bar{F}\otimes h^{\otimes 2n}\otimes F)\right)\right]^{1/2n}\\ &\quad=\frac{1}{n}\left(\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n+2k)}q^{(n+k)-\operatorname{\mathrm{cyc}}_{0}(\pi(\alpha^{-1}\cup(01\ldots n)\cup\alpha))}C_{\pi}(\bar{F}\otimes h^{\otimes 2n}\otimes F)\right)^{1/2n}\\ &\quad\leq\frac{1}{n}\left(\frac{(2n+2k)!}{2^{n+k}(n+k)!}\left\|F\right\|^{2}\left\|h\right\|^{2n}\right)^{1/2n}\\ &\quad\sim\frac{1}{n}2^{1/2}(n+k)^{1/2}e^{-1/2}\left\|h\right\|\rightarrow 0,\end{split}$
where we used the fact that $\left|q\right|\leq 1$ and $\left|\operatorname{\mathcal{P}}_{2}(2n)\right|=\frac{(2n!)}{2^{n}n!}$. ∎
@@
$\begin{split}a^{-}_{(01)}(h)(\alpha\otimes F)&=q\sum_{k\neq\alpha^{-1}(0)}P^{[0,n]\setminus\left\{k\right\}}_{[0,n-1]}((0k)\alpha)|_{\left\{k\right\}^{c}}\otimes r^{-}_{k}(h)F\\ &\qquad+\delta_{\alpha(0)\neq 0}P^{[0,n]\setminus\left\{\alpha^{-1}(0)\right\}}_{[0,n-1]}\alpha|_{\left\{\alpha^{-1}(0)\right\}^{c}}\otimes r^{-}_{\alpha^{-1}(0)}(h)F.\end{split}$
The distribution of $\mathnormal{W}\left({(01)}\otimes h\right)$ is the unnormalized average empirical distribution of a GUE matrix with mean 0 and variance $\left\|h\right\|$.
@@
$\begin{split}&\mathnormal{W}\left(\alpha\otimes F\right)\ \mathnormal{W}\left((01)\otimes h\right)\\ &\quad=\mathnormal{W}\left((0\ n+1)\alpha\otimes r^{+}(h)F\right)\\ &\quad\quad+\sum_{k=1}^{n}q^{\operatorname{\mathrm{cyc}}_{0}(((0k)\alpha)|_{\left\{k\right\}^{c}})-\operatorname{\mathrm{cyc}}_{0}((0k)\alpha)+1}\mathnormal{W}\left(P^{[0,n]\setminus\left\{k\right\}}_{[0,n-1]}((0k)\alpha)|_{\left\{k\right\}^{c}}\otimes r^{-}_{k}(h)F\right)\\ &\quad=\mathnormal{W}\left((0\ n+1)\alpha\otimes r^{+}(h)F\right)\\ &\quad\quad+q\sum_{k\neq\alpha^{-1}(0)}\mathnormal{W}\left(P^{[0,n]\setminus\left\{k\right\}}_{[0,n-1]}((0k)\alpha)|_{\left\{k\right\}^{c}}\otimes r^{-}_{k}(h)F\right)\\ &\quad\quad+\delta_{\alpha(0)\neq 0}\mathnormal{W}\left(P^{[0,n]\setminus\left\{\alpha^{-1}(0)\right\}}_{[0,n-1]}\alpha|_{\left\{\alpha^{-1}(0)\right\}^{c}}\otimes r^{-}_{\alpha^{-1}(0)}(h)F\right).\end{split}$
@@
which should be compared with Theorem 22.12 in [32]. ∎
@@ 5.2. The relation to a construction by Bożejko and Guţă
We contrast the algebra $\mathcal{P}(\mathcal{H}_{\mathbb{R}})$ with a construction from [6]. In section 5 of that paper, Bożejko and Guţă considered the Fock space with the inner product
@@
and the corresponding expression is zero for an odd number of factors (compare with equation (13)). Here $c(\pi)$ is again the number of cycles of a permutation corresponding to a partition $\pi$, but this correspondence is more subtle. For a pair partition $\pi$, there is a unique non-crossing partition $\tilde{\pi}$ with the same openers and closers as $\pi$. If $i\stackrel{{\scriptstyle\pi}}{{\sim}}j$ and $i\stackrel{{\scriptstyle\tilde{\pi}}}{{\sim}}k$, then for the corresponding permutation $\sigma$, $\sigma(i)=j$ if $i<j$, and $\sigma(i)=k$ if $k<i$ (it is easy to check that this is an alternative). In other words, $\sigma$ is a permutation with an upper partition $\pi$ and the lower partition $\tilde{\pi}$ in the sense of Corteel [9]. Then $c(\pi)$ is the number of cycles of $\sigma$.
@@
Both our construction and the one in [6] are quite different from the more familiar $q$-deformed free Fock space, for which $q=\pm\frac{1}{N}$ hold no special meaning.
Remark 5.38.
Combining our construction with [6], one could consider a Fock space with the inner product
@@
Here the coefficient of $\beta$ is the generating function of the number of primitive factorizations of $\beta$. See [31].
@@ 6.1. Background
Abstract trace polynomials, and in particular the expression $\operatorname{\mathrm{tr}}_{\alpha}[x_{1},\ldots,x_{n}]$ for $\alpha\in S_{0}(n)$, were defined in the introduction. Let $\mathcal{A}$ be a unital algebra, $\mathcal{C}$ its center, and $F:\mathcal{A}\rightarrow\mathcal{C}$ a unital, tracial, $\mathcal{C}$-bimodule linear map. For any $\alpha\in S_{0}(n)$, we can similarly form $F_{\alpha}(a_{1},\ldots,a_{n})\in\mathcal{A}$, and consider it as the application of the trace monomial $\operatorname{\mathrm{tr}}_{\alpha}[x_{1},\ldots,x_{n}]$ to the elements $a_{1},\ldots,a_{n}$. See [8]. We extend the notation to $F_{\eta}$ for $\eta\in\mathbb{C}[S_{0}(n)]$ by linearity.
@@ 6.1.1. Invariant theory of $N\times N$ matrices
where $\operatorname{\mathrm{Tr}}$ is the (un-normalized) trace on $M_{N}(\mathbb{C})$. Indeed, since for the purposes of this expansion, $x_{ij}^{(k)}$ and $(x_{ij}^{(k)})^{\ast}$ can be considered as independent variables, this follows directly from the first Procesi-Razmyslov theorem [33, 36]. The result is usually formulated using $GL(n)$-invariance. However, the argument ultimately reduces to Schur-Weyl duality, for which (in the case of inner product spaces) unitary invariance is sufficient.
@@ 6.1.2. Hermitian Brownian motion
Let $\left\{b_{ij}(h):h\in\mathcal{H}_{\mathbb{R}}\right\}$ be $N^{2}$ standard Gaussian processes indexed by the same real Hilbert space as in Section 2.4, independent for different $(i,j)$, represented on the same probability space. Define the $N\times N$ Hermitian Gaussian process $\left\{X(h):h\in\mathcal{H}_{\mathbb{R}}\right\}$ by
[…]
@@
The following Proposition is a slight extension of a well-known result (see for example Lemma 22.31 of [32]) to the case of a general permutation $\alpha\in S_{0}(n)$ and GUE matrices indexed by a Hilbert space. We provide the proof for completeness.
@@
We now restrict the index set $S$ to be a real Hilbert space $\mathcal{H}_{\mathbb{R}}$. Denote
@@ 6.2. The isomorphisms
Numerous corollaries follow by combining Theorem 6.4 with results from earlier in the article. We only list a few of them explicitly. Others include Proposition 4.9 (chaos decomposition in the univariate case), 5.7 (expansion of the Hermite polynomial and stochastic integral), 5.11 (product formula for Hermite polynomials and stochastic integrals), 5.17 (linearization coefficients), and Corollary 5.12 (Factorization of the Hermite polynomial for orthogonal arguments).
@@
(Compare with Proposition 5.19.) Let $\mathcal{H}=L^{2}(\mathbb{R}_{+},dx)$. Then for each $\alpha$, the process $H_{\alpha}(X(\mathbf{1}_{[0,t]}))$ is a martingale.
For Hermite polynomials of matrix argument, the martingale property was proved in [29] by generating function methods. See Section 6.2.1 for the connection.
@@
In the case $\mathcal{H}_{\mathbb{R}}=L^{2}(\mathbb{R}_{+},dx)$ and $F\in L^{2}(\mathbb{R}_{+}^{n},dx^{\otimes n})$, we may identify the random matrix $\mathcal{E}[\mathnormal{W}\left(\eta\otimes_{s}F\right)]$ with a stochastic integral
@@
$\mathcal{E}[\mathnormal{W}\left(\eta\otimes(\mathbf{1}_{J_{1}}\otimes\ldots\otimes\mathbf{1}_{J_{n}})\right)]=\operatorname{\mathrm{Tr}}_{\eta}[X(\mathbf{1}_{J_{1}}),\ldots,X(\mathbf{1}_{J_{n}})]\\$
@@
Let $\mathcal{H}_{\mathbb{R}}$ be a real Hilbert space, $\mathcal{K}_{\mathbb{R}}=M_{N}(\mathbb{R})\otimes\mathcal{H}_{\mathbb{R}}$, and $\mathcal{K}=M_{N}(\mathbb{C})\otimes\mathcal{H}_{\mathbb{R}}$ its complexification. Let $Y\in\mathcal{A}^{\mathrm{equiv}}_{N}(\mathcal{K})$, with the inner product induced by the (Gaussian) expectation. For a polynomial variable $X$ in the Gaussian Hilbert space indexed by $\mathcal{H}_{\mathbb{R}}$, recall the notation :X::\!\!X\!\!: from Section 2.4. Extend this operation to $Y$ entry-wise:
$(:\!\!Y\!\!:)_{ij}=:\!\!(Y_{ij})\!\!:.$
Then :Y::\!\!Y\!\!: is also equivariant.
@@
It suffices to prove the result when each $Y_{ij}$ is a homogeneous polynomial of degree $n$. Let $Y=P(X(h):h\in\mathcal{K})$ and $Y\!\!:=Q(X(h):h\in\mathcal{K})$. Let $U\in U_{N}(\mathbb{C})$. Since each $(Y_{ij}):$ is orthogonal to $\mathcal{P}_{n-1}(\mathcal{K})$, for any such polynomial $R$,
@@
The matrix with entries $(\operatorname{\mathrm{Tr}}_{\alpha}[X(h_{1}),\ldots,X(h_{n})])_{\ell r}\!\!:$ has leading term $\operatorname{\mathrm{Tr}}_{\alpha}[X(h_{1}),\ldots,X(h_{n})]$, polynomial entries, and (by Lemma 6.10) is equivariant. Therefore it is of the form
@@
In [3], Biane defined the matricial Segal-Bargmann transform. We outline the construction, and send the reader to the original article for details. The ordinary Segal-Bargmann transform is defined in Section 2.4. Apply that construction with $\mathcal{H}_{\mathbb{R}}=M_{N}^{sa}(\mathbb{C})$, the space of Hermitian matrices. Upgrade the Segal-Bargmann transform to its matricial version by applying it entry-by-entry. Then it follows from Theorem 6.11 that, as in Lemma 7 and Theorem 7 of [3],
@@
(18) $h_{\lambda}(x_{1},\ldots,x_{N})=\mathcal{E}[\mathnormal{W}\left(\lambda\otimes h^{\otimes n}\right)].$
@@
For $\lambda\in\operatorname{\mathrm{Par}}(n)$, the Hermite polynomial of matrix argument (for $\beta=2$) is the symmetric polynomial in $\left\{x_{1},\ldots,x_{N}\right\}$ with leading term $\frac{\left|\lambda\right|!}{c_{\lambda}}s_{\lambda}$ which is an eigenfunction of the operator $D^{\ast}-E^{\ast}$ with eigenvalue n-n (note a misprint in [13]). Here (see Corollary 7.1.7.4 in [37])
@@
is the Schur polynomial and $c_{\lambda}=\prod(\lambda_{i}+\lambda_{j}^{\prime}-i-j+1)$ is the hook length. See [2, 13, 14] for more details.